A Uniform chain with mass per unit length L is lifted upwards with a constant speed by a force F. Find out the force in terms of time
1
Expert's answer
2020-12-10T11:07:27-0500

Chain has a constant speed. Therefore, net force on it should be zero.


"F=F_t+W"


"F_t=\frac{dm}{dt}(v_0-0)=\frac{dm}{dl}\frac{dl}{dt}v_0=\rho v_0^2"


"W=(y\cdot\rho)\cdot g"


"F=\rho v_0^2+\rho gy=\rho(v_0^2+gy)=\rho(v_0^2+g\cdot (v_0t))=\rho v_0(v_0+gt)" . Answer